OpenStudy (anonymous):

what is log small 6 25 to four decimal points?

OpenStudy (luigi0210):

Plug it into a calculator? Or use the change in base formula.

OpenStudy (anonymous):

\[\log _{6}25\]

OpenStudy (jdoe0001):

\(\bf \textit{change of base formula of }log_ab=\cfrac{log_cb}{log_ca}\qquad thus \\ \quad \\ log_625\implies \cfrac{log_{10}25}{log_{10}6} \)

OpenStudy (campbell_st):

well a simple way to handle it is the change to base to base 10 or base e I'll show you base e \[\log_{6}(25) = \frac{\ln(25)}{\ln(6)}\] now you can just pick up a scientific calculator and evaluate

OpenStudy (anonymous):

yeah im not good with calculators..care to further elaborate the equation?

OpenStudy (jdoe0001):

http://desmos.com/calculator

OpenStudy (campbell_st):

well just enter \[\ln(25) \div \ln(6) = \] that will give you an answer and here is a calculator it works on http://web2.0calc.com/

OpenStudy (anonymous):

thanks.